# The SNO+ Collaboration

### v2

@@ VII Results and Conclusions
Figure 10: The neutron yield measured by different experiments with varying average muon energies. The black empty markers represent measurements made by other experiments [an_cosmogenic_2018, aglietta_neutron_1989, zhao_measurement_2022, abe_production_2010, boehm_neutron_2000, hertenberger_muon-induced_1995, blyth_measurement_2016, bellini_cosmogenic_2013, collaboration_measurement_1999, collaboration_measurement_2023, aharmim_cosmogenic_2019]. The solid black line shows the fit made by Daya Bay using some of the liquid scintillator results [an_cosmogenic_2018]. The blue round marker is the SNO+ result of this paper, and the blue square marker shows the GEANT4 predicted result for SNO+. The dashed lines represent the FLUKA-based predictions for liquid scintillator: Wang et al. [wang_predicting_2001] (yellow) and Kudryavtsey et al. [kudryavtsev_simulations_2003] (green).
The GEANT4 prediction for the neutron yield in water at SNO+ has been calculated to be $Y_{n}=(2.130\pm 0.001(\textup{stat)})\times 10^{-4}\,\textup{cm}^{2}\textup{g}^{-1}\mu^{-1}$. This prediction for the neutron yield is $\sim 30\,\%$ less than the result measured. A similar deficit in neutron production between FLUKA and GEANT4 of up to $30\,\%$ was observed for muons of energy $\gtrsim 100\,\textup{GeV}$ in hydrocarbon materials [araujo_muon-induced_2005]. Daya Bay also reported a similar observation between the two models in liquid scintillator [an_cosmogenic_2018]. We note that the discrepancy appears to be almost entirely due to single neutron events; as can be seen in Figure 4, the GEANT4 prediction shows an excess in muons producing only one neutron compared with data. This effect was also observed in [an_cosmogenic_2018].
@@ Acknowledgements.
Capital funds for SNO+ were provided by the Canada Foundation for Innovation and matching partners: Ontario Ministry of Research, Innovation and Science, Alberta Science and Research Investments Program, Queen’s University at Kingston, and the Federal Economic Development Agency for Northern Ontario. This research was supported by (Canada) the Natural Sciences and Engineering Research Council of Canada, the Canadian Institute for Advanced Research, the Ontario Early Researcher Awards; (U.S.) the Department of Energy (DOE) Office of Nuclear Physics, the National Science Foundation, and the DOE National Nuclear Security Administration through the Nuclear Science and Security Consortium; (U.K.) the Science and Technology Facilities Council, the Royal Society, and an NMES studentship from King’s College London; (Portugal) Fundação para a Ciência e a Tecnologia (FCT-Portugal); (Germany) the Deutsche Forschungsgemeinschaft; (Mexico) DGAPA-UNAM and Consejo Nacional de Ciencia y Tecnología; and (China) the Discipline Construction Fund of Shandong University. We also thank SNOLAB and SNO+ technical staff; the Digital Research Alliance of Canada; the GridPP Collaboration and support from Rutherford Appleton Laboratory, the King’s College London Computational Research, Engineering and Technology Environment (CREATE), and the Savio computational cluster at the University of California, Berkeley. Additional long-term storage was provided by the Fermilab Scientific Computing Division. For the purposes of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising. Representations of the data relevant to the conclusions drawn here are provided within this paper.

### v3

@@ VII Results and Conclusions
Figure 10: The neutron yield measured by different experiments with varying average muon energies. The black empty markers represent measurements made by other experiments [3, 10, 12, 7, 5, 4, 6, 9, 8, 13, 11]. The solid black line shows the fit made by Daya Bay using some of the liquid scintillator results [3]. The blue round marker is the SNO+ result of this paper, and the blue square marker shows the GEANT4 predicted result for SNO+. The dashed lines represent the FLUKA-based predictions for liquid scintillator: Wang et al. [1] (yellow) and Kudryavtsey et al. [32] (green).
The GEANT4 prediction for the neutron yield in water at SNO+ has been calculated to be $Y_{n}=(2.130\pm 0.001(\textup{stat)})\times 10^{-4}\,\textup{cm}^{2}\textup{g}^{-1}\mu^{-1}$. This prediction for the neutron yield is $\sim 30\,\%$ less than the result measured. A similar deficit in neutron production between FLUKA and GEANT4 of up to $30\,\%$ was observed for muons of energy $\gtrsim 100\,\textup{GeV}$ in hydrocarbon materials [33]. Daya Bay also reported a similar observation between the two models in liquid scintillator [3]. We note that the discrepancy appears to be almost entirely due to single neutron events; as can be seen in Figure 4, the GEANT4 prediction shows an excess in muons producing only one neutron compared with data. This effect was also observed in [3].
@@ Acknowledgements.
Capital funds for SNO+ were provided by the Canada Foundation for Innovation and matching partners: Ontario Ministry of Research, Innovation and Science, Alberta Science and Research Investments Program, Queen’s University at Kingston, and the Federal Economic Development Agency for Northern Ontario. This research was supported by (Canada) the Natural Sciences and Engineering Research Council of Canada, the Canadian Institute for Advanced Research, the Ontario Early Researcher Awards; (U.S.) the Department of Energy (DOE) Office of Nuclear Physics, the National Science Foundation, and the DOE National Nuclear Security Administration through the Nuclear Science and Security Consortium; (U.K.) the Science and Technology Facilities Council, the Royal Society, and an NMES studentship from King’s College London; (Portugal) Fundação para a Ciência e a Tecnologia (FCT-Portugal); (Germany) the Deutsche Forschungsgemeinschaft; (Mexico) DGAPA-UNAM and Consejo Nacional de Ciencia y Tecnología; and (China) the Discipline Construction Fund of Shandong University. We also thank SNOLAB and SNO+ technical staff; the Digital Research Alliance of Canada; the GridPP Collaboration and support from Rutherford Appleton Laboratory, the King’s College London Computational Research, Engineering and Technology Environment (CREATE), and the Savio computational cluster at the University of California, Berkeley. Additional long-term storage was provided by the Fermilab Scientific Computing Division. For the purposes of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising. Representations of the data relevant to the conclusions drawn here are provided within this paper and its supplemental material [34].